OpenStudy (ohmyquaffle):

The slope-intercept form of the equation of a line that passes through point (–3, 8) is y = –x + 6. What is the point-slope form of the equation for this line?

pooja195 (pooja195):

Hi there! Welcome to OpenStudy! :~) Any ideas on where to start?

OpenStudy (ohmyquaffle):

not really haha

pooja195 (pooja195):

Alright that's fine :) Point slope formula: \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1) Can you tell me the following: x1=? y1=? m=?

OpenStudy (ohmyquaffle):

x1= -3 y1=8 m=-1?

pooja195 (pooja195):

Perfect now just plug them into the slope formula what do you get?

OpenStudy (ohmyquaffle):

thank you!

pooja195 (pooja195):

You're welcome :)